# TianQuan-S2S: A Subseasonal-to-Seasonal Global Weather Model via Incorporate Climatology State

### v4

@@ Abstract
Accurate Subseasonal-to-Seasonal (S2S) forecasting is vital for decision-making in agriculture, energy production, and emergency management. However, it remains a challenging and underexplored problem due to the chaotic nature of the weather system. Recent data-driven studies have shown promising results, but their performance is limited by the inadequate incorporation of climate states and a model tendency to degrade, progressively losing fine-scale details and yielding over-smoothed forecasts. To overcome these limitations, we propose TianQuan-S2S, a global S2S forecasting model that integrates initial weather states with climatological means via incorporating climatology into patch embedding and enhancing variability capture through an uncertainty-augmented Transformer. Extensive experiments on the Earth Reanalysis 5 (ERA5) reanalysis dataset demonstrate that our model yields a significant improvement in both deterministic and ensemble forecasting over the climatology mean, traditional numerical methods, and data-driven models. Ablation studies empirically show the effectiveness of our model designs. Remarkably, our model outperforms skillful numerical ECMWF-S2S and advanced data-driven Fuxi-S2S in key meteorological variables.
@@ 1 Introduction
However, data-driven subseasonal forecasting remains challenging and less explored  (Chen et al., 2024; Liu et al., 2025). The extended time horizon renders the initial weather conditions and key variables inadequate for making accurate predictions due to the chaotic nature of the weather system (Mariotti et al., 2018). Deterministic machine learning models are typically trained on data from recent historical data, which provides limited support for subseasonal forecasting (He et al., 2019). As the forecast horizon extends, the initial state becomes less informative, resulting in blurred and unrealistic predictions. This limitation hampers the effectiveness of machine learning models on subseasonal time scales, despite their importance for climate-related applications (Lam et al., 2023; Bonavita, 2023).
@@ 2 Preliminaries
In weather forecasting, perturbed forecasting is used to generate ensemble predictions by adding small perturbations to the initial atmospheric state (Zhou et al., 2022). Because the atmosphere is a chaotic system, even minor uncertainties in the initial conditions can lead to large differences in forecast outcomes (Nakashita & Enomoto, 2025). Perturbed forecasting generates multiple forecasts from slightly modified initial states to capture possible atmospheric evolutions. In addition to perturbing initial conditions, ensemble systems also introduce perturbations within the model itself. For example, ECMWF and NCEP apply stochastic parameterization schemes such as the Stochastically Perturbed Parametrizations (SPP), which randomly modify model tendencies like convection or turbulence (Tsiringakis et al., 2024). Moreover, data-driven methods such as SwinVRNN use learned distribution perturbations to generate diverse ensemble members, improving forecast reliability by explicitly modeling uncertainty (Hu et al., 2023).
@@
$\bm{\hat{X}}_{t_{15}:t_{45}}=f_{\Theta}(\bm{G},\bm{X}_{clim},\bm{X}_{t_{-4}:t_{1}}),$ (1)
where $\Theta$ denotes the parameters of neural networks. $\bm{\hat{X}}_{t_{20}:t_{45}}$ is the predicted value from day 15 to day 45. In this work, we employ the 5-day historical data $\bm{X}_{t_{-4}:t_{1}}$ as the input, which we empirically found beneficial for model performance.
@@ 3.1 Patch Embedding
Previously, some researchers (Zhang & Patel, 2018; Wang et al., 2021; Kumar & Tiwari, 2023) have utilized edge priors to enhance details. Building on this work, we introduce a multi-layer convolutional layer to extract climate information $\bm{F_{clim}}\in\mathbb{R}^{H\times W\times K}$ from the climatology $\bm{X}_{clim}$. In this convolution, pixel pair differences are calculated, and prior information is explicitly encoded into the CNN, enhancing the model’s representation by learning valuable trend information.
@@
Compute the spatial weights as $\bm{W}_{att}=\mathrm{unreshape(}\sigma\!\left(\frac{\bm{Q}^{(v)}\bm{K}^{(v)}}{\sqrt{D}}\right)\bm{V}^{(v)})\in[0,1]^{H\times W\times C}$. Finally, a convolutional layer maps the fused representation to the output $\bm{F}\in\mathbb{R}^{H\times W\times K}$:
@@
where 1 denotes a matrix with all elements equal to 1. In summary, the attention-based fusion module adaptively balances and integrates complementary features, ensuring more effective fusion for long-term climate prediction.
@@
where [·] denotes C channel-wise concatenation and * denotes that broadcasting is applied during the computing. Subsequently, the initial embedding $\bm{E}$ is fed into the Transformer for further processing.
@@ 3.2 Transformer
The Vision Transformer (ViT) architecture has proven effective for various prediction tasks, including climate modeling. By processing input data through stacked attention blocks, the transformer captures both local and global dependencies within the data. In the context of weather forecasting, each block learns hierarchical representations.
@@
To improve generalization and handle forecast uncertainty, we add an uncertainty block in each transformer block. This block injects Gaussian noise into the network, allowing it to learn data distributions in all lead times. Specifically, an uncertainty block introduces Gaussian noise at each transformer block to produce the next prediction $\bm{E}^{(n+1)}$, as described by the following formula:
@@ 4 Experiments
We conducted experiments using the ERA5 dataset (Hersbach et al., 2020), covering 40 years (1979-2018) with global hourly data across multiple pressure levels and the surface. To improve the accuracy of long-term predictions, we preprocess the hourly ERA5 data by averaging it daily at six-hour intervals. Bilinear interpolation is used to downsample the data to 5.625° (32 × 64 grid points) and 1.40625° (128 × 256 grid points), while selecting the 13 most critical pressure levels. For more details on data processing, refer to Appendix D.4. We obtained the dataset variables as shown in Table 10 in Appendix D. We divided the daily averaged ERA5 dataset into training (1979-2015), validation (2016), and testing (2017-2018) sets based on years.
@@
Following existing works (Bi et al., 2023; Rasp et al., 2024), we primarily quantify the performance improvement of the model using latitude-weighted RMSE and Anomaly Correlation Coefficient (ACC), and then assess its long-term forecasting capability through Continuous Ranked Probability Score (CRPS), Spread Mean Error (SME), and Relative Quantile Error (RQE). All of these metrics can be found in the Appendix E.
@@
We run our approach once to generate a single time value for deterministic forecasting. In addition, as an AI-based method, our model can perform large-member ensemble forecasts with small computational costs. Following the ensemble strategy adopted in FuXi-S2S (Chen et al., 2024), we constructed ensemble forecasting for ClimaX (Nguyen et al., 2023) and TianQuan-S2S by introducing random perturbations to the initial weather states. Specifically, we generated 50 random perturbations and added them to the unperturbed initial state $x_{t}$. Thus, we obtain a 51-member ensemble forecast and compute an ensemble mean forecast by simply averaging the forecast results.
@@
T2m
FuXi-S2S 2.450 2.454 2.484 2.520 2.542 2.563 -
Wind10
FuXi-S2S 3.407 3.574 3.668 3.678 3.771 3.752 -
@@
We used the AdamW optimizer (Kingma, 2014; Loshchilov & Hutter, 2017) with parameters $\lambda_{1}$ = 0.9 and $\lambda_{2}$ = 0.99. A weight decay of $1e^{-5}$ was applied to all parameters except the positional embeddings. The learning rate was set to $5e^{-5}$, with a linear warmup schedule over 5000 steps (5 epochs), followed by a cosine annealing schedule over 95000 steps (95 epochs). Our model training was implemented using PyTorch (Paszke et al., 2019). In the training process, we train multiple models with different lead time settings, where the lead time ranges from 15 to 45 days with an interval of 5 days. The output from each model, corresponding to a single step prediction, is combined to form a 45-day subseasonal forecast target. The model was trained on eight GPUs with 80GB of memory, achieving 77.6 TFLOPS of computing power.
@@ 4.1 Main Results
  • Regardless of whether the resolution is 1.40625° or 5.625°, TianQuan-Cliamte consistently outperforms all baselines in all variables. Specifically, the average RMSE improves by 0.14 on T850 ($K$), 59 on Z500 ($m^{2}/s^{2}$), and 0.353 on Wind10 ($m/s$) compared to the best baseline, all representing significant improvements.
  • Compared with the transformer-based direct prediction model ClimaX, FuXi-S2S (00 member) performs worse on T850, Z500, and Wind10 before 25 days due to iterative error accumulation at each step. However, beyond 25 days, FuXi-S2S stabilizes and avoids the model collapse and severe metric degradation observed in ClimaX, while our TianQuan-S2S model remains even more stable.
@@
To provide a global view of model predictions, we visualize the day 35 prediction distribution in Figure 4. As shown in the figure, all models show a global trend of change, but the results of ClimaX have a serious loss of details. FuXi-S2S shows high value changes on both T2m and T850, but its distribution on Wind10 is significantly different from the ground truth. In contrast, our model aligns more closely with the realistic observations provided by ERA5, easing the model collapse problem. In addition, to analyze the correlation between predictions and the ground truth, we draw day 25 scatter plot in Figure 5. It demonstrates that our method is closer to the true values, and outperforms numerical model ECMWF-S2S and data-driven models, including FuXi-S2S and ClimaX across correlation ($R^{2}$), mean average error (MAE), and Bias metrics. After comparing these details, we further validated our framework of incorporating climatology state by attention-based fusion. More visualizations, such as different variable predictions, can be found in the Appendix 22.
@@ 4.2 Ablation Study
To validate the effect of each model design on the overall model performance, we compare four experimental models based on whether they use attention-based fusion as well as noise generation. The results are shown in Table 3, where we can obtain the following observations that:
@@ 5 Relative Work
Weather forecasting mainly involves gridded prediction tasks, akin to image-to-image translation. Recent advancements have been made in short- to medium-term weather and climate forecasting. Notable examples include FourCastNet (Pathak et al., 2022), which was the first to attempt multi-variable forecasting in the short term with results comparable to NWP models. Pangu-Weather (Bi et al., 2023) and GraphCast (Lam et al., 2022) surpass traditional numerical methods in achieving multi-variable forecasting in the medium term. Based on the framework of transformer, FuXi (Chen et al., 2023b) and FengWu (Chen et al., 2023a) developed a model method adapted to the Earth system. NeuralGCM (Kochkov et al., 2024) is designed based on the GCM architecture to capture spatial dependencies in weather prediction. Exploring higher resolution forecasting, FengWu-GHR (Han et al., 2024) aims to improve prediction accuracy. Aurora (Bodnar et al., 2025) and Prithvi-WxC (Schmude et al., 2024) both focus on Weather forecasting based on foundation models. Aardvark Weather (Allen et al., 2025) and FuXi Weather (Sun et al., 2025) both employ end-to-end deep learning models, directly processing raw observational data and assimilating satellite information for weather forecasting. While these advancements extend forecasting capabilities, the absence of effective climate information leads to a significant decline in performance. This highlights the need for new approaches specifically designed for subseasonal forecasting.
@@
Subseasonal forecasting, which provides predictions 2 to 6 weeks ahead, fills the crucial gap between short-term weather forecasts (typically up to 15 days) and long-term climate predictions that extend to seasonal and longer timescales. Although machine learning models have achieved significant progress in medium-range and climate prediction, their effectiveness in subseasonal forecasting has been less pronounced (Wang et al., 2024; He et al., 2021). However, at sub-seasonal to seasonal time scales, the predictive results of machine learning methods often lack the details of forecasting (Vitart et al., 2022; Nathaniel et al., 2024). This limitation leads to iterative errors in long-term forecasts, ultimately rendering the results unusable (Lam et al., 2023). Recent methods focus on building more effective subseasonal forecasting models to enhance predictive skill (Chen et al., 2024; Liu et al., 2025). To strengthen subseasonal forecasting capability, we proposed a patch embedding that incorporates climatological information and an uncertainty-augmented Transformer that captures weather variability. In particular, although both Fuxi-S2S and TianQuan-S2S add perturbation in model, Variational Autoencoder-based FuXi-S2S adds perturbations only within the latent space, while TianQuan-S2S injects learnable Gaussian noise directly into the feature representations at every layer of Transformer blocks, continuously reinforces variability, and more effectively mitigates model collapse in long-lead forecasts. In addition, Fuxi-S2S does not explicitly incorporate climate information.
6 Conclusion

### v5

@@ Abstract
Accurate Subseasonal-to-Seasonal (S2S) forecasting is vital for decision-making in agriculture, energy production, and emergency management. However, it remains a challenging and underexplored problem due to the chaotic nature of the weather system. Recent data-driven studies have shown promising results, but their performance is limited by the inadequate incorporation of climate states and a model tendency to degrade, progressively losing fine-scale details and yielding over-smoothed forecasts. To overcome these limitations, we propose TianQuan-S2S, a global S2S forecasting model that integrates initial weather states with climatological means via incorporating climatology into patch embedding and enhancing variability capture through an uncertainty-augmented Transformer. Extensive experiments on the Earth Reanalysis 5 (ERA5) reanalysis dataset demonstrate that our model yields a significant improvement in both deterministic and ensemble forecasting over the climatology mean, traditional numerical methods, and data-driven models. Ablation studies empirically show the effectiveness of our model designs. Remarkably, our model outperforms skillful numerical ECMWF-S2S and advanced data-driven Fuxi-S2S in key meteorological variables. The code implementation can be found in https://github.com/zhangminglang42/TianQuan.
22footnotetext: Corresponding authors
@@ 1 Introduction
However, data-driven subseasonal forecasting remains challenging and less explored  (Chen et al., 2024; Liu et al., 2025a). The extended time horizon renders the initial weather conditions and key variables inadequate for making accurate predictions due to the chaotic nature of the weather system (Mariotti et al., 2018). Deterministic machine learning models are typically trained on data from recent historical data, which provides limited support for subseasonal forecasting (He et al., 2019). As the forecast horizon extends, the initial state becomes less informative, resulting in blurred and unrealistic predictions. This limitation hampers the effectiveness of machine learning models on subseasonal time scales, despite their importance for climate-related applications (Lam et al., 2023).
@@ 2 Preliminaries
In weather forecasting, perturbed forecasting is used to generate ensemble predictions by adding small perturbations to the initial atmospheric state (Zhou et al., 2022). Because the atmosphere is a chaotic system, even minor uncertainties in the initial conditions can lead to large differences in forecast outcomes (Nakashita and Enomoto, 2025). Perturbed forecasting generates multiple forecasts from slightly modified initial states to capture possible atmospheric evolutions. In addition to perturbing initial conditions, ensemble systems also introduce perturbations within the model itself. For example, ECMWF and NCEP apply stochastic parameterization schemes such as the Stochastically Perturbed Parametrizations (SPP), which randomly modify model tendencies like convection or turbulence (Tsiringakis et al., 2024). Moreover, data-driven methods such as SwinVRNN use learned distribution perturbations to generate diverse ensemble members, improving forecast reliability by explicitly modeling uncertainty (Hu et al., 2023).
@@
$\bm{\hat{X}_{t_{15}:t_{45}}}=f_{\Theta}(\bm{G},\bm{X}_{clim},\bm{X}_{t_{-4}:t_{0}}),$ (1)
where $\Theta$ denotes the parameters of neural networks. $\bm{\hat{X}}_{t_{15}:t_{45}}$ is the predicted value from day 15 to day 45. In this work, we employ the 5-day historical data $\bm{X}_{t_{-4}:t_{0}}$ as the input, which we empirically found beneficial for model performance.
@@ 3.1 Patch Embedding
Previously, some researchers (Zhang and Patel, 2018; Wang et al., 2021; Kumar and Tiwari, 2023) have utilized edge priors to enhance details. Building on this work, we introduce a multi-layer convolutional layer to extract climate information $\bm{F_{clim}}\in\mathbb{R}^{H\times W\times K}$ from the climatology $\bm{X}_{clim}$. In this convolution, pixel pair differences are calculated, and prior information is explicitly encoded into the CNN, enhancing the model’s representation by learning valuable trend information.
@@
Compute the spatial weights as $\bm{W}_{att}=\mathrm{unreshape(}\sigma\!\left(\frac{\bm{Q}^{(v)}\bm{K}^{(v)}}{\sqrt{N}}\right)\bm{V}^{(v)})\in[0,1]^{H\times W\times C}$. Finally, a convolutional layer maps the fused representation to the output $\bm{F}\in\mathbb{R}^{H\times W\times K}$:
@@
where 1 denotes a matrix with all elements equal to 1. In summary, the attention-based fusion adaptively integrates complementary features, ensuring more effective fusion for long-term prediction.
@@
where [·] denotes C channel-wise concatenation and * denotes that broadcasting is applied during the computation. Subsequently, this embedding $\bm{E}$ is fed into the Transformer for further processing.
@@ 3.2 Transformer
The Vision Transformer (ViT) architecture has proven effective for various prediction tasks, including climate modeling (Nguyen et al., 2023). By processing input data with stacked attention blocks, the transformer captures both local and global dependencies. In the context of weather forecasting, each block learns hierarchical representations.
@@
In general, adding stochastic perturbations helps prevent the forecast from collapsing onto a single deterministic trajectory and better captures the growth of uncertainty at extended lead times. In particular, instead of perturbing only the initial conditions, which lose impact as lead time increases (Buizza et al., 2005), we inject learnable Gaussian perturbations at each Transformer layer during the forward pass, stabilizing forecasts and mitigating over-smoothing at extended ranges.
Specifically, an uncertainty block introduces Gaussian noise at each transformer block to produce the next prediction $\bm{E}^{(n+1)}$, as described by the following formula:
@@ 4 Experiments
We conducted experiments using the ERA5 dataset (Hersbach et al., 2020), covering 40 years (1979-2018) with global hourly data across multiple pressure levels and the surface. To improve the accuracy of long-term predictions, we preprocess the hourly ERA5 data by averaging it daily at six-hour intervals. Bilinear interpolation is used to downsample the data to 5.625° (32 × 64 grid points) and 1.40625° (128 × 256 grid points), while selecting the 13 most critical pressure levels. For more details on data processing, refer to Appendix F.4. We obtained the dataset variables as shown in Table 10 in Appendix F. We divided the daily averaged ERA5 dataset into training (1979-2015), validation (2016), and testing (2017-2018) sets based on years.
@@
Following existing works (Bi et al., 2023; Rasp et al., 2024), we primarily quantify the performance improvement of the model using latitude-weighted RMSE and Anomaly Correlation Coefficient (ACC), and then assess its long-term forecasting capability through Continuous Ranked Probability Score (CRPS), Spread Mean Error (SME), and Relative Quantile Error (RQE). All of these metrics can be found in the Appendix G.
@@
We run our approach once to generate a single time value for deterministic forecasting. In addition, as an AI-based method, our model can perform large-member ensemble forecasts with small computational costs. For ClimaX, we implemented ensemble forecasting by introducing random perturbations to the initial weather states. In the case of TianQuan-S2S, we incorporated learnable noise at each block level to generate multiple forecasts. Specifically, we created 50 random perturbations, which were then added to the original unperturbed state. This approach results in a 51-member ensemble, and the ensemble mean forecast is computed by averaging the individual forecast outputs.
@@
T2m FuXi-S2S 2.450 2.454 2.484 2.520 2.542 2.563 - Wind10 FuXi-S2S 3.407 3.574 3.668 3.678 3.771 3.752 -
@@
We used the AdamW optimizer (Kingma, 2014; Loshchilov and Hutter, 2017) with parameters $\lambda_{1}$ = 0.9 and $\lambda_{2}$ = 0.99. A weight decay of $1e^{-5}$ was applied to all parameters except the positional embeddings. The learning rate was set to $5e^{-5}$, with a linear warmup schedule over 5000 steps (5 epochs), followed by a cosine annealing schedule over 95000 steps (95 epochs). Our model training was implemented using PyTorch (Paszke et al., 2019). In the training process, we train multiple models with different lead time settings, where the lead time ranges from 15 to 45 days with an interval of 5 days. The output from each model, corresponding to a single step prediction, is combined to form a 45-day subseasonal forecast target. The model was trained on eight GPUs with 80GB of memory, achieving 77.6 TFLOPS of computing power.
@@ 4.1 Main Results
  • Regardless of whether the resolution is 1.40625° or 5.625°, TianQuan-S2S consistently outperforms all baselines in all variables. Specifically, the average RMSE improves by 0.14 on T850 ($K$), 59 on Z500 ($m^{2}/s^{2}$), and 0.353 on Wind10 ($m/s$) compared to the best baseline, all representing significant improvements.
  • Compared with the transformer-based direct prediction model ClimaX, FuXi-S2S (00 member) performs worse on T850, Z500, and Wind10 before 25 days due to iterative error accumulation at each step. However, beyond 25 days, FuXi-S2S stabilizes and avoids the model collapse and severe metric degradation observed in ClimaX. Unlike ClimaX, TianQuan-S2S adds attention-based climatology fusion and learnable Gaussian noise in the Transformer blocks to anchor forecasts, reduce drift, and stabilize long-lead predictions.
@@
  • We note that for Wind10 at S2S lead times, both FuXi-S2S and our model are worse than climatology, mainly because ML-based ensemble means over-smooth this highly variable field and thus increase RMSE. In contrast, the 61-day sliding-window climatology used in WeatherBench2 (Rasp et al., 2024) yields a much smoother and more stable Wind10 baseline, especially at long leads.
@@
To provide a global view of model predictions, we visualize the day 30 prediction distribution in Figure 4. As shown in the figure, all models show a global trend of change, but the results of ClimaX have a serious loss of details. FuXi-S2S shows high value changes on both T2m and T850, but its distribution on Wind10 is significantly different from the ground truth. In contrast, our model aligns more closely with the realistic observations provided by ERA5, easing the model collapse problem. In addition, to analyze the correlation between predictions and the ground truth, we draw day 25 scatter plot in Figure 5. It demonstrates that our method is closer to the true values, and outperforms numerical model ECMWF-S2S and data-driven models, including FuXi-S2S and ClimaX, across correlation ($R^{2}$), mean average error (MAE), and Bias metrics. After comparing these details, we further validated our framework of incorporating climatology state by attention-based fusion. More visualizations, such as different variable predictions, can be found in the Appendix 22.
@@ 4.2 Ablation Study
Effectiveness of the Model Design
To validate the effect of each model design on the overall model performance, we compare four experimental models based on whether they use attention-based fusion as well as noise injection, as summarized in Table 3. The main observations are:
@@
Table 4: Comparison of Fusion and Perturbation Methods in Ensemble Mean RMSE. Our proposed method (Default) achieves the best performance compared to other fusion strategies, perturbation methods, and varying the number of noise injection layers.
Methods 15 20 25 30 35 40 45 Methods 15 20 25 30 35 40 45
T2m Concat 2.525 2.556 2.565 2.583 2.627 2.655 2.707 Z500 Concat 830 814 808 837 845 819 832
Gate 2.476 2.528 2.530 2.541 2.563 2.589 2.670 Gate 811 802 797 818 828 802 818
IC Perturb 2.446 2.506 2.527 2.538 2.556 2.624 2.674 IC Perturb 814 812 809 821 823 808 818
FLN 2.497 2.556 2.560 2.583 2.601 2.616 2.617 FLN 825 814 818 827 825 816 823
Layer 1 Only 2.599 2.626 2.653 2.714 2.777 2.802 2.841 Layer 1 Only 817 811 806 818 823 807 821
Layer 1-2 2.530 2.562 2.606 2.623 2.693 2.699 2.712 Layer 1-2 810 809 798 816 818 792 811
Layer 1-4 2.535 2.548 2.593 2.603 2.593 2.652 2.729 Layer 1-4 814 808 802 810 816 785 817
Layer 1-6 2.511 2.578 2.602 2.567 2.623 2.657 2.707 Layer 1-6 807 797 772 811 812 779 815
Layer 1-7 2.456 2.496 2.518 2.532 2.554 2.604 2.624 Layer 1-7 799 782 776 804 813 778 808
[…]
Analysis of Fusion and Perturbation Methods
We investigate our method’s superiority in three areas: (1) Fusion strategies, comparing Concatenation (i.e., Concat, merging climatology on channels) and a Learnable Gate (i.e., adaptive climatology selection per pixel); (2) Perturbation methods, comparing Initial-Condition Perturbation (i.e., IC Perturb, perturbing only initial conditions) and Fixed-Layers Noise (i.e., FLN, removing learnable functions); (3) Impact of noise injection layers, where ”Layers 1-2” refers to injecting noise into the 1st and 2nd blocks. Results in Table 4 show that:
  • Comparing the two variants of fusion strategies and ensemble forecast perturbations on T2m and Z500 at all lead times, our method (Default) achieves the best performance, confirming that attention is a more effective and interpretable way to exploit climatological priors, and indicating that the proposed learnable noise injection is a superior alternative to Initial-Condition Perturbations and better emulates the climate evolution process than Fixed-Layer Noise.
  • Investigating the impact of uncertainty blocks, we further conduct an ablation study on per-layer perturbation. As the number of perturbed layers increases, the overall performance consistently improves: the ensemble-mean RMSE of T2m and Z500 is reduced by up to 0.223 and 28.85, respectively, demonstrating that noise injection enhances forecast accuracy.
Figure 6: RMSE comparison between different Gaussian noise scales at all lead times. A lighter color indicates better results: overall, the effect is better when $\sigma=1$.
Impact of Noise Scale
To further understand the impact of the standard deviation of Gaussian noise, we compared different noise standard deviations $\sigma$ and obtained the results shown in Figure 6. From the results, we can find that when $\sigma$ is around 1, the perturbation model generates better results, indicating that this noise scale strikes an optimal balance between regularization and model accuracy.
@@ 5 Relative Work
Weather forecasting mainly involves gridded prediction tasks, akin to image-to-image translation. Recent advancements (Zheng et al., 2025) have been made in short- to medium-term weather and climate forecasting. Notable examples include FourCastNet (Pathak et al., 2022), which was the first to attempt multi-variable forecasting in the short term with results comparable to NWP models. Pangu-Weather (Bi et al., 2023) and GraphCast (Lam et al., 2022) surpass traditional numerical methods in achieving multi-variable forecasting in the medium term. Based on the framework of transformer, FuXi (Chen et al., 2023b) and FengWu (Chen et al., 2023a) developed a model method adapted to the Earth system. NeuralGCM (Kochkov et al., 2024) is designed based on the GCM architecture to capture spatial dependencies in weather prediction. Exploring higher resolution forecasting, FengWu-GHR (Han et al., 2024) aims to improve prediction accuracy. Aurora (Bodnar et al., 2025) and Prithvi-WxC (Schmude et al., 2024) both focus on Weather forecasting based on foundation models. Aardvark Weather (Allen et al., 2025) and FuXi Weather (Sun et al., 2025) both employ end-to-end deep learning models, directly processing raw observational data and assimilating satellite information for weather forecasting. While these advancements extend forecasting capabilities, the absence of effective climate information leads to a significant decline in performance. This highlights the need for new approaches specifically designed for subseasonal forecasting.
@@
Subseasonal forecasting, which provides predictions 2 to 6 weeks ahead, fills the crucial gap between short-term weather forecasts (typically up to 15 days) and long-term climate predictions that extend to seasonal and longer timescales. Although machine learning models have achieved significant progress in medium-range and climate prediction, their effectiveness in subseasonal forecasting has been less pronounced (Wang et al., 2024; He et al., 2021). However, at sub-seasonal to seasonal time scales, the predictive results of machine learning methods often lack the details of forecasting (Vitart et al., 2022; Nathaniel et al., 2024). This limitation leads to iterative errors in long-term forecasts, ultimately rendering the results unusable (Lam et al., 2023). Recent methods focus on building more effective subseasonal forecasting models to enhance predictive skill (Chen et al., 2024; Liu et al., 2025a; b). To strengthen subseasonal forecasting capability, we proposed a patch embedding that incorporates climatological information and an uncertainty-augmented Transformer that captures weather variability. In particular, although both Fuxi-S2S and TianQuan-S2S add perturbations in the model, Variational Autoencoder-based FuXi-S2S adds perturbations only within the latent space, while TianQuan-S2S injects learnable Gaussian noise directly into the feature representations at every layer of Transformer blocks, continuously reinforces variability, and more effectively mitigates model collapse in long-lead forecasts. In addition, Fuxi-S2S does not explicitly incorporate climate information.
6 Conclusion and Future Work